6 2. Design an active filter to achieve the following transfer function: V. (0 Vi 3 + 2x10²s + 106
Answer #1

Griven Vo = 106 s²+2x1035 + 100 Now compairing in with standard and order trans for function given by Wa? 5²+2Ewns twn2 Wn =Now and order Actuie filter wing Opamp is given as: GH Vinom R3 Av R4 (2 R2 RI Open loop 12 here Av is a voltage gain It R2 =by compairing with standard equation we get RC= I wo RC = 1 RCE 103 #RC= 10-3 Let us take R= loka i.e lotr 3 こ 10 10-7E 104 *